Ram is three times as old as his son.After ten year the sum of their ages will be 76 years .find their present ag
Answer by Edwin McCravy(20056)

Ram is three times as old as his son.

Ram's age = 3 times son's age

Let son's age = x

Then Ram's age = 3x

After ten years,...

After 10 years son will be x+10
After 10 years Ram will be 3x+10

the sum of their ages will be 76 years .

That says:

(x+10)+(3x+10) = 76
    x+10+3x+10 = 76
         4x+20 = 76
            4x = 56
             x = 14    
So 

Son is 14 

Ram is 3 times 14 or 42
 
Edwin